OpenStudy (star):

Determine whether the vectors {(2, 0, 1) , (3, 1, 2) , (1, 1, 1) , (7, 3, 5)} span R^3 ? I'm really stuck on how to find the "span" of a matrix or anything.

OpenStudy (anonymous):

create a matrix using those vectors as columns, and row reduce it.

OpenStudy (star):

i reduced it to \[=\left[\begin{matrix}2 & 0 & 1 \\ 0 & -6 & -3\end{matrix}\right]\] but apparently it is supposed to be (2,0,1) and (1,1,1)...?

OpenStudy (anonymous):

Once you row reduce you need to look for the pivots, the columns the pivots are in represent the columns of the original matrix that are linearly independent.

OpenStudy (anonymous):

So the only vectors that are linearly independent are (2,0,1) and (3,1,2). In fact: (1,1,1)= (3,1,2)-(2,0,1) and (7,3,5)=3*(3,1,2)-(2,0,1) the last two vectors are just linear combinations of the first two. They dont contribute any no information. Since only 2 vectors are linearly independent, the span is 2 dimensional. its a plane.

OpenStudy (anonymous):

new*, not no. my bad.

OpenStudy (star):

oh! thank you so much i understand it now! thank youuu